Q: The field winding of a d.c. motor is connected directly across a 440 V supply. When the winding is at the room temperature of 17 °C, the current is 2.3 A. After the machine has been running for some hours, the current has fallen to 1.9 A, the voltage rem aining unaltered. Calculate the average temperature throughout the winding, assuming the temperature coefficient of resistance of copper to be 0.004 26/°C at 0 °C.
Answer #1

The resistance of the winding at room temperature (170C ), R17 = ( V / I1) = ( 440 / 2.3) = 191.304 ohms

Let us assume that the average temperature of the winding is t 0C .

Rt = ( V / I2 ) = ( 440 / 1.9) = 231.579 ohms

Let us assume the resistance of the winding at 0 0 C is R0

The temperature coefficient of resistance of Cu,